The length of a rectangle is 5 times the width and its area is 180 square units. What is the length of the rectangle?
Found 2 solutions by nerdybill, cleomenius:
Answer by nerdybill(7384)

The length of a rectangle is 5 times the width and its area is 180 square units. What is the length of the rectangle?
.
Let w = width
then
5w = length
.
w(5w) = 180
5w^2 = 180
w^2 = 36
w = 6 units (width)
.
Length:
5w = 5(6) = 30 units

Answer by cleomenius(959)

width = x
height = 5x
Area = 5x^2 = 180
x^2 = 36
x = 6
6 units is the width,.
30 units is the height.
6 * 30 = 180 units squared.
Cleomneius